40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 4 19:01:59 2022, Tega, Update, Computers, Front-end machine in supermicro boxes IMG_20220804_184444473.jpgIMG_20220804_191658206.jpgIMG_20220804_185336240.jpgIMG_20220804_185023002.jpg
    Reply  Mon Aug 8 17:16:51 2022, Tega, Update, Computers, Front-end machine setup IMG_20220808_170349717.jpg
       Reply  Wed Aug 10 20:51:14 2022, Tega, Update, Computers, CDS upgrade Front-end machine setup IMG_20220810_171002928.jpgIMG_20220810_171019633.jpg
          Reply  Tue Aug 16 18:22:59 2022, Tega, Update, Computers, c1teststand rack mounting for CDS upgrade IMG_20220816_180157132.jpgIMG_20220816_175125874.jpg
             Reply  Wed Aug 17 11:10:51 2022, rana, Update, Computers, c1teststand rack mounting for CDS upgrade 
             Reply  Mon Aug 22 19:02:15 2022, Tega, Update, Computers, c1teststand rack mounting for CDS upgrade II IMG_20220822_185437763.jpgIMG_20220822_131340214.jpgc1bhd.jpegc1lsc.jpeg
                Reply  Tue Aug 23 22:30:24 2022, Tega, Update, Computers, c1teststand OS upgrade - I 
                Reply  Fri Aug 26 14:05:09 2022, Tega, Update, Computers, rack reshuffle proposal for CDS upgrade 6x
                   Reply  Sun Aug 28 23:14:22 2022, Jamie, Update, Computers, rack reshuffle proposal for CDS upgrade 
                      Reply  Mon Aug 29 15:15:46 2022, Tega, Update, Computers, 3 FEs from LLO got delivered today IMG_20220829_145533452.jpgIMG_20220829_144801365.jpg
                         Reply  Tue Aug 30 15:21:27 2022, Tega, Update, Computers, 3 FEs from LHO got delivered today IMG_20220830_144925325.jpgIMG_20220830_142307495.jpgIMG_20220830_143059443.jpg
                   Reply  Mon Sep 19 20:21:06 2022, Tega, Update, Computers, 1X7 and 1X6 work IMG_20220919_204013819.jpgIMG_20220919_203541114.jpgIMG_20220919_203458952.jpg
                      Reply  Tue Sep 20 23:06:23 2022, Tega, Update, Computers, Setup the 6 new front-ends to boot off the FB1 clone IMG_20220921_084220465.jpgdolphin_err_init_state.pngdolphin_final_state.png
                         Reply  Wed Sep 21 17:16:14 2022, Tega, Update, Computers, Setup the 6 new front-ends to boot off the FB1 clone 
                            Reply  Thu Sep 22 20:57:16 2022, Tega, Update, Computers, build, install and start 40m models on teststand  dolphin_state_plus_c1shimmer.pngFE_status_overview.png
                               Reply  Fri Sep 23 19:07:03 2022, Tega, Update, Computers, Work to improve stability of 40m models running on teststand  dolphin_state_all_green.pngdolphin_state_IPC_glitch.png
                            Reply  Thu Sep 29 15:12:02 2022, JC, Update, Computers, Setup the 6 new front-ends to boot off the FB1 clone 
                               Reply  Tue Oct 4 21:00:49 2022, Chris, Update, Computers, Failed takeover attempt with the new front ends 
                                  Reply  Thu Oct 6 07:29:30 2022, Chris, Update, Computers, Successful takeover attempt with the new front ends 
Message ID: 17164     Entry time: Thu Sep 29 15:12:02 2022     In reply to: 17151     Reply to this: 17172
Author: JC 
Type: Update 
Category: Computers 
Subject: Setup the 6 new front-ends to boot off the FB1 clone 

[Jamie, Christopher, JC]

This morning we decided to label the the fiber optic cables. While doing this, we noticed that the ends had different label, 'Host' and 'Target'. Come to find out, the fiber optic cables are directional. Four out of Six of the cables were reversed. Luckily, 1 cable for the 1Y3 IO Chassis has a spare already laid (The cable we are currently using).  Chris, Jamie, and I have begun reversing these cable to there correct position.

Quote:

[Tega, JC]

We laid 4 out of 6 fiber cables today. The remaining 2 cables are for the I/O chassis on the vertex so we would test the cables the lay it tomorrow. We were also able to identify the problems with the 2 supposedly faulty cable, which are not faulty. One of them had a small bend in the connector that I was able to straighten out with a small plier and the other was a loose connection in the switch part. So there was no faulty cable, which is great! Chris wrote a matlab script that does the migration of all the model files. I am going through them, i.e. looking at the CDS parameter block to check that all is well. Next task is to build and install the updated models. Also need to update the '/opt/rtcds' and '/opt/rtapps' directory to the latest in the 40m chiara system.

 

 

ELOG V3.1.3-